# FPC-VLA: A Vision-Language-Action Framework with a Supervisor for Failure Prediction and Correction

### v1

@@ IV-E Ablation Study
Table VI presents the ablation study on FPC-VLA, focusing on the Supervisor and Action Fusion Module. “w/o” indicates “without”. Results are measured on four WidowX tasks. Removing the Supervisor greatly degrades performance, as it corrects deviations from the expert trajectory. Removing corrections during data generation and inference also lowers success, indicating that VLM’s correction suggestions aid action optimization. Providing in-context examples further improves success by clarifying task requirements and output format. For the Action Fusion Module, similarity-guided fusion outperforms both no fusion and simple averaging by better leveraging multi-modal actions, while the time-decay strategy boosts success by emphasizing recent predictions.
@@
To further evaluate the Supervisor’s effect on robustness, we perform an action perturbation experiment with and without the Supervisor. The end-effector pose is randomly disturbed within the range of [0.001, 0.01]. Results show that although perturbations reduce success rates in both settings, the drop is 39.3% without the Supervisor but only 20.9% with FPC-VLA.
V Conclusion
@@
In this paper, we propose FPC-VLA, a VLA Framework with a supervisor for failure prediction and correction. By integrating a VLM-based supervisor and a similarity-guided action fusion module, FPC-VLA not only refines robot actions before execution but also proactively mitigates potential failures. Extensive experiments across simulated environments and real-world robotic platforms demonstrate that FPC-VLA achieves superior performance over existing methods in both zero-shot and fine-tuned settings, highlighting its robustness, generalization ability, and potential for safe and reliable deployment in open-ended robotic applications.

### v2

@@ 4.6 Ablation Study
Table 7 presents the ablation study on FPC-VLA, focusing on the Supervisor and Action Fusion Module. “w/o” indicates “without”. Results are measured on WidowX, Xiaomi robot and ALOHA tasks. Removing the Supervisor greatly degrades performance, as it corrects deviations from the expert trajectory. Removing corrections during data generation and inference also lowers success, indicating that VLM’s correction suggestions aid action optimization. Providing in-context examples further improves success by clarifying task requirements and output format. For the Action Fusion Module, similarity-guided fusion outperforms both no fusion and simple averaging by better leveraging multi-modal actions, while the time-decay strategy boosts success by emphasizing recent predictions.
@@
To further evaluate the Supervisor’s effect on robustness, we perform an action perturbation experiment with and without the Supervisor. The end-effector pose is randomly disturbed within the range of [0.01, 0.1]. Results show that although perturbations reduce success rates in both settings, the drop is 31.3% without the Supervisor but only 16% with FPC-VLA.
5 Conclusion
@@
In this paper, we propose FPC-VLA, a VLA Framework with a supervisor for failure prediction and correction. By integrating a VLM-based supervisor and a dual-stream action fusion module, FPC-VLA not only refines robot actions before execution but also proactively mitigates potential failures. Extensive experiments across simulated environments and real-world robotic platforms demonstrate that FPC-VLA achieves superior performance over existing methods in zero-shot and fine-tuned settings, highlighting its robustness, generalization ability, and potential for safe and reliable deployment in open-ended robotic applications.
6 Limitations and Future Work
While FPC-VLA demonstrates strong performance in failure prediction and correction, it has some limitations. First, the supervisor is triggered exclusively by gripper state changes, making it unable to address post-grasp failures—such as objects slipping or dropping after successful pickup—which are common in real-world manipulation. Mitigating such issues typically requires force/torque sensing and compliant control strategies, yet current VLA frameworks, including ours, rely solely on visual inputs and lack the necessary modalities to perceive or react to contact forces. Integrating force sensors into the VLA pipeline could enable more robust handling of dynamic interactions. Second, although our experiments include long-horizon tasks (e.g., drawer opening and stove operation), their complexity and duration remain limited compared to the demands of real-world domestic service scenarios, which often involve hours of continuous, multi-step reasoning under uncertainty. Nevertheless, as VLA architectures continue to evolve—with improvements in memory, planning, and multimodal grounding—we believe this paradigm holds significant promise for scaling to truly complex, everyday robotic applications.